Question 10 Pendant Companys break-even point in sales is $690,000 and its variable expenses are 60% of sales. If the company had a profit of $10,000 in 2014, its sales must have been- $762,000 $665,000 $700,000 。$715,000
Answer #1

Contribution margin=Sales-Variable costs

=(100-60)=40%

At breakeven point;Contribution margin=Fixed cost=($690,000*0.4)=$276000

Hence for target profits;Target Contribution margin=Target profits+Fixed costs

=(10,000+276,000)=$286,000

Hence target sales=$286,000/Contribution margin ratio

=$286,000/0.4

=$715,000